Z Instructions Question 1 10 pts Two industrial robots are being considered for automation purposes. Using a present worth analysis at 10% interest per year, determine the best alternative Robot X Robot Y 95,000 5000 in year 1 increasing by $1,000 per year 4,000 First Cost S$ 5,000 Annual Cost$/year 9,000 Salvage value, Useful Life, years Upload Choose a Filte 20 3
Answer #1

First of all take the LCM of useful life in order to do Present worth analysis

LCM of 2 and 4 is 4. So, we have to do present worth analysis for 4 years.

Robot X

Given, initial cost, I = $ 55,000

Annual cost, A = $ 9,000

Salvage value, S = $ 0

Time, n = 4

PWx =-1-A(P/A, i%, n)-(1-S)(P/ F, i%,n)+S(P/F, i%, n)

PWx =-55,000-9000(P/A,10%,4)-55,000(P/F,10%,2)+0

PWx = -55,000-9,000*3.170 -55,000*0.8264

PWx = - $ 128,982

Now do the analysis for Robot Y

PWy = -95,000 - A(P/A,10%,4)-G(P/G,10%,4) +S(P/F,10%,4)

PWy =-95,000-5,000*3.170-1,000*4.378 + 4,000*0.6830

PWy = -$ 112,496

Select Robot Y.

Please contact if having any query. Thank you!